ADNAN MUHAMMED

Grade 12,

solve the eq |x/x-1|+|x|=x^2/|x-1|

solve the eq |x/x-1|+|x|=x^2/|x-1|

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

|x/x-1|+|x|=x^2/|x-1|

|x| + |x||x-1|=x2

condition x can be equal to 1 because x-1 is on denominator

case 1 x<0

-x +x2-x  =x2

 x=0   which is not  a solution

case 2      0≤x<1

 x -(x)(x-1) =x

 x(x-1)=0

 x=0 and 1

 x=1 is not a solution

case 3    1<x

x+x(x-1) =x2

x=x2  which is always true 

so final solution          x=0   and x>1


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin